OpenStudy (anonymous):

Find the exact area of a circle having the given circumference. 3pi

OpenStudy (anonymous):

Circumference = 2*pi*radius = 3*pi so radius = 3/2 Area = pi*r^2 = 9/4*pi

OpenStudy (anonymous):

C = 2*pi*r c = 3*pi r = 3/2 A = pi*r^2 A = pi*(3/2)*(3/2) A = 9/4pi